OpenStudy (anonymous):

what is the equation of the line that passes through (0, 4) and has a slope of -3

OpenStudy (luigi0210):

Use the point-slope formula \(\large y-y_{1}=m(x-x_{1})\)

OpenStudy (texaschic101):

or just use y = mx + b y = mx + b slope(m) = -3 (0,4) x = 0 and y = 4 now sub 4 = -3(0) + b 4 = 0 + b 4 = b so your equation is : y = -3x + 4